# Shannon Capacity of Channels with Markov Insertions, Deletions and Substitutions

### v1

@@ 3 Functions of channels and their properties
This proposition can be interpreted as follows. The process of forming the (random) output of channel $W$ can be seen as a pipeline of two steps. First, choose the value of $z$ based on probabilities $W^{\prime}(z|x)$. This is equivalent to channel $W^{\prime}$. Second, randomly choose one value among all $\leq 2^{\Phi}$ possible values of $y\in f^{-1}(y^{\prime})$. The maximum entropy of the second step is achieved when all preimages are equiprobable, namely, the entropy of the second step is at most $\Phi$.
For a channel sequence $U_{n}:{\mathcal{X}}^{n}\rightsquigarrow{\mathcal{Y}}_{n}$ and sequence of functions $f_{n}:{\mathcal{Y}}_{n}\to{\mathcal{Z}}_{n}$, we denote by $V={f}{\left({U}\right)}$ a channel sequence $V_{n}={f_{n}}{\left({U_{n}}\right)}:{\mathcal{X}}^{n}\rightsquigarrow{\mathcal{Z}}_{n}$, $n\in{\mathbb{N}}$. Also, we denote by $\Phi_{n}=\log_{2}\max_{z\in{\mathcal{Z}}_{n}}\left|{f_{n}^{-1}(z)}\right|$.
@@
If $\Phi_{n}\in o(n)$, then ${\mathbf{C}}(U)={\mathbf{C}}(V)$. If, furthermore, at least one of ${\mathbf{I}}(U)$, ${\mathbf{I}}(V)$ exist, then they both exist and ${\mathbf{I}}(U)={\mathbf{I}}(V)$.
@@
First, note that ${\mathbf{E}}_{p_{n}\circ V_{n}}{\left[{\mathbf{i}(V_{n},p_{n},x,y)}\right]}={\mathbf{E}}_{p_{n}\circ U_{n}}{\left[{\mathbf{i}(V_{n},p_{n},x,f(y))}\right]}$. Using Proposition 2, for any $p_{n}$:
${\mathbf{I}}(U_{n},p_{n})-{\mathbf{I}}(V_{n},p_{n})={\mathbf{E}}_{p_{n}\circ U_{n}}{\left[{\mathbf{i}(U_{n},p_{n},x,y)-\mathbf{i}(V_{n},p_{n},x,f_{n}(y))}\right]}$
$\leq{\mathbf{E}}_{p_{n}\circ U_{n}}{\left[{\left|{\mathbf{i}(U_{n},p_{n},x,y)-\mathbf{i}(V_{n},p_{n},x,f_{n}(y))}\right|}\right]}\leq\Phi_{n}.$
In particular, for any sequence of input distributions $\left\{{p_{n}\in{\mathbb{D}}_{{\mathcal{X}}^{n}}}\right\}_{n\in{\mathbb{N}}}$,
$0\displaystyle 0\leq\lim_{n\to\infty}{\mathbf{E}}_{p_{n}\circ U_{n}}{\left[{\left|{\frac{\mathbf{i}(U_{n},p_{n},x,y)-\mathbf{i}(V_{n},p_{n},x,f_{n}(y))}{n}}\right|}\right]}\ \leq\lim_{n\to\infty}\frac{\Phi_{n}}{n}=$ 0 . (13)
@@
$\forall\varepsilon>0,\delta>0,\exists n^{*}_{0}:\forall n\geq n_{0},\exists p^{*}:~{}\Pr_{p^{*}\circ U_{n}(x,y)}{\left[{\frac{\mathbf{i}(U_{n},p^{*},x,y)}{n}>\alpha-\varepsilon}\right]}$ $>1-\delta.$ (14)
$\exists\varepsilon^{*}>0,\delta^{*}>0:\forall n_{0},\exists n^{*}\geq n_{0}:\forall p,~{}\Pr_{p\circ U_{n^{*}}(x,y)}{\left[{\frac{\mathbf{i}(V_{n^{*}},p,x,f_{n^{*}}(y))}{n^{*}}\leq\alpha-\varepsilon^{*}}\right]}$ $\geq\delta^{*}.$ (15)
@@
$\exists\varepsilon^{*}>0,\delta^{*}>0:\begin{cases}\exists n^{*}_{0}:\forall n\geq n^{*}_{0},\exists p_{n}^{*}:~{}\Pr_{p_{n}^{*}\circ U_{n}(x,y)}{\left[{\mathbf{i}(U_{n},p_{n}^{*},x,y)/n>\alpha-\varepsilon^{*}/2}\right]}>1-\delta^{*}/2\\ \forall n_{0},\exists n^{*}\geq n_{0}:\forall p_{n^{*}},~{}\Pr_{p_{n^{*}}\circ U_{n^{*}}(x,y)}{\left[{\mathbf{i}(V_{n^{*}},p_{n^{*}},x,f_{n^{*}}(y))/{n^{*}}\leq\alpha-\varepsilon^{*}}\right]}\geq\delta^{*}\end{cases}.$
So, the first inequality holds for all $n$ starting from $n^{*}_{0}$, and for each $n$ specific $p_{n}^{*}$ should be picked up. The second inequality holds for some $n$ larger than arbitrary $n_{0}$, for all $p_{n}$. Bringing these two conditions together, we can satisfy both inequalities by doing as follows:
@@
  • since $n^{*}\geq n^{*}_{0}$, the first equality still holds for some specific value of $p^{*}_{n^{*}}$. The second inequality will hold for $p_{n^{*}}^{*}$ as well, since it holds for any $p_{n^{*}}$.
@@
$*.\displaystyle\exists\varepsilon^{*}>0,\delta^{*}>0:\forall n_{0},\exists n\geq n_{0},\exists p_{n}:\begin{cases}\Pr_{p_{n}\circ U_{n}(x,y)}{\left[{\mathbf{i}(U_{n},p_{n},x,y)/n>\alpha-\varepsilon^{*}/2}\right]}>1-\delta^{*}/2\\ \Pr_{p_{n}\circ U_{n}(x,y)}{\left[{\mathbf{i}(V_{n},p_{n},x,f_{n}(y))/{n}\leq\alpha-\varepsilon^{*}}\right]}\geq\delta^{*}\end{cases}.$
Note that both probabilities are defined over the same probability space. The sum of the two probabilities is greater than $21+\delta^{*}/$ 2, which means that with probability greater than $\delta^{*}/2$ the difference $\mathbf{i}(U_{n},p_{n},x,y)/n-\mathbf{i}(V_{n},p_{n},x,f_{n}(y))/{n}$ is greater than $\varepsilon^{*}/2$, which implies
$\exists\varepsilon^{*}>0,\delta^{*}>0:\forall n_{0},\exists n\geq n_{0},\exists p_{n}:{\mathbf{E}}_{p_{n}\circ U_{n}}{\left[{\left|{\frac{\mathbf{i}(U_{n},p_{n},x,y)-\mathbf{i}(V_{n},p_{n},x,f_{n}(y))}{n}}\right|}\right]}>\frac{\varepsilon^{*}\delta^{*}}{4},$
which contradicts (13).
Equality of i-capacities follows from (13). Let $p^{*}_{n}$ be the optimal distribution for $U_{n}$. If either ${\mathbf{I}}(U)$ or ${\mathbf{I}}(V)$ exists, then
$0\leq{\mathbf{I}}(U)-{\mathbf{I}}(V)\leq\lim_{n\to\infty}\frac{{\mathbf{I}}(U_{n},p^{*}_{n})-{\mathbf{I}}(V_{n},p^{*}_{n})}{n}\leq\lim_{n\to\infty}\frac{\Phi_{n}}{n}=0,$
so ${\mathbf{I}}(U)={\mathbf{I}}(V)$. ∎
Thus, one can apply such function sequences to a channel sequence $W$, changing neither ${\mathbf{I}}(W)$ nor ${\mathbf{C}}(W)$.
@@
Consider channels $U:{\mathcal{X}}\rightsquigarrow{\mathcal{Y}}$ and $V:{\mathcal{X}}\rightsquigarrow{\mathcal{Z}}$, such that $V=f(U)$, $f:{\mathcal{Y}}\to{\mathcal{Z}}$. Consider partitions ${\mathcal{Y}}={\mathcal{A}}\sqcup{\mathcal{B}}$, ${\mathcal{Z}}={\mathcal{A}}^{\prime}\sqcup{\mathcal{B}}^{\prime}$, such that $f({\mathcal{A}})={\mathcal{A}}^{\prime}$ and $f$ is bijective between ${\mathcal{A}}$ and ${\mathcal{A}}^{\prime}$. Denote by $\beta(x)=\Pr_{U(y|x)}{\left[{y\in{\mathcal{B}}}\right]}$ the probability that the output of $U$ belongs to ${\mathcal{B}}$. If $\forall x:\beta(x)\leq\overline{\beta}$, then
${\mathbf{I}}(U)-{\mathbf{I}}(V)\leq\overline{\beta}\log_{2}|{\mathcal{X}}|+\frac{1}{e\ln 2}.$ (16)
@@
Consider some $p\in{\mathbb{D}}_{\mathcal{X}}$. By definition, ${\mathbf{I}}(U,p)={\mathbf{E}}_{p\circ U(x,y)}{\left[{\mathbf{i}(U,p,x,y)}\right]}$. The summation, corresponding to taking the expectation, can be expressed by independent summation over ${\mathcal{A}}$ and ${\mathcal{B}}$:
@@
Let us analyze term $b$. First, introduce modified distribution ${\widetilde{p}}(x)=p(x)\beta(x)/\Delta$, where $\Delta=\sum_{x}p(x)\beta(x)$. Note that $0\leq\Delta\leq\overline{\beta}$. Substituting $p(x)U(y|x)=\Delta{\widetilde{p}}(x)U^{\prime}(y|x)$ for $y\in{\mathcal{B}}$, one obtains
@@
$=\Delta\cdot{\mathbf{I}}(U^{\prime},{\widetilde{p}})+\Delta\sum_{y\in{\mathcal{B}},x\in{\mathcal{X}}}{\widetilde{p}}(x)U^{\prime}(y|x)\log_{2}\frac{\beta(x)}{\Delta}$
$=\Delta{\left({{\mathbf{I}}(U^{\prime},{\widetilde{p}})-\log_{2}\Delta}\right)}+\sum_{y\in{\mathcal{B}},x\in{\mathcal{X}}}p(x)U^{\prime}(y|x)\beta(x)\log_{2}\beta(x)$
@@
Consider channel sequences $U_{n}:{\mathcal{X}}^{n}\rightsquigarrow{\mathcal{Y}}_{n}$ and $V_{n}:{\mathcal{X}}^{n}\rightsquigarrow{\mathcal{Z}}_{n}$, such that $V_{n}=f_{n}(U_{n})$, $f_{n}:{\mathcal{Y}}_{n}\to{\mathcal{Z}}_{n}$. Consider partitions ${\mathcal{Y}}_{n}={\mathcal{A}}_{n}\sqcup{\mathcal{B}}_{n}$, ${\mathcal{Z}}_{n}={\mathcal{A}}^{\prime}_{n}\sqcup{\mathcal{B}}^{\prime}_{n}$, such that $f_{n}({\mathcal{A}}_{n})={\mathcal{A}}^{\prime}_{n}$ and $f_{n}$ is bijective between ${\mathcal{A}}_{n}$ and ${\mathcal{A}}^{\prime}_{n}$. Denote by $\beta_{n}(x)=\Pr_{U_{n}(y|x)}{\left[{y\in{\mathcal{B}}_{n}}\right]}$, and $\overline{\beta}_{n}=\max_{x}\beta_{n}(x)$. If the output of channel $U_{n}$ almost always belongs to ${\mathcal{A}}_{n}$, i.e., $\lim_{n\to\infty}\overline{\beta}_{n}=0,$ then ${\mathbf{C}}(U)={\mathbf{C}}(V)$. If, in addition, one of ${\mathbf{I}}(U)$, ${\mathbf{I}}(V)$ exists, then they both exist and ${\mathbf{I}}(U)={\mathbf{I}}(V)$.
@@
If either ${\mathbf{I}}(U)$ or ${\mathbf{I}}(V)$ exists, then, using proposition 4,
${\mathbf{I}}(U)-{\mathbf{I}}(V)=\lim_{n\to\infty}\frac{{\mathbf{I}}(U_{n})-{\mathbf{I}}(V_{n})}{n}=\lim_{n\to\infty}{\left({\frac{\overline{\beta}_{n}n\log_{2}|{\mathcal{X}}|}{n}+\frac{1}{ne\ln 2}}\right)}=0,$
and both ${\mathbf{I}}(U)$ and ${\mathbf{I}}(V)$ exist.
According to (9), proving that ${\mathbf{C}}(U)={\mathbf{C}}(V)$ can be done analyzing the probability of $\mathbf{i}(V_{n},p_{n},x,y)$ being less than $\alpha n$ for some value of $\alpha$. If $\mathbf{i}(V_{n},p_{n},x,y)=\mathbf{i}(U_{n},p_{n},x,y)$ and $\mathbf{i}(U_{n},p_{n},x,y)>\alpha n$, then surely $\mathbf{i}(V_{n},p_{n},x,y)>\alpha n$. This can be used as a lower bound on probability:
@@
$\leq\Pr_{p_{n}\circ U_{n}(x,y)}{\left[{\mathbf{i}(V_{n},p_{n},x,f(y))\neq\mathbf{i}(U_{n},p_{n},x,y)}\right]}+\Pr_{p_{n}\circ U_{n}(x,y)}{\left[{\mathbf{i}(U_{n},p_{n},x,y)\leq\alpha n}\right]}$
In the last expression, with $n\to\infty$, the first probability goes to zero for any $p_{n}$. So,
@@
which means that the set of such $\alpha$’s for $U_{n}$ is a subset of such set for $V_{n}$, which implies ${\mathbf{C}}(V)\geq{\mathbf{C}}(U)$ and thus ${\mathbf{C}}(V)={\mathbf{C}}(U)$.
If this proposition is combined with Proposition 3, the result is that one can apply any finite number of functions, which either have sufficiently small pre-image or sufficiently small probability of changing the output, without changing neither ${\mathbf{I}}$ nor ${\mathbf{C}}$ of the original channel sequence.
4 The Markov-IDS channel sequence
4.1 The IDS channel sequence
In [3] a general channel model, which represents any channel with independent insertions, deletions, and substitutions is considered.
Consider a basic channel $W:{\mathcal{X}}\rightsquigarrow{\overline{\mathcal{Y}}}$, such that for each $x\in{\mathcal{X}}$ the expected length ${\mathbf{N}}(y)$ of the output sequence $y\in{\overline{\mathcal{Y}}}$ is finite:

### v3

@@ III Functions of channels and their properties
This proposition can be interpreted as follows. The process of forming the (random) output of channel $W$ can be seen as a pipeline of two steps. First, choose the value of $z$ based on probabilities $W^{\prime}(z|x)$. This is equivalent to channel $W^{\prime}$. Second, randomly choose one value among all (at most) $2^{\Phi}$ possible values of $y\in f^{-1}(y^{\prime})$. The maximum entropy of the second step is achieved when all preimages are equiprobable, namely, the entropy of the second step is at most $\Phi$.
For a channel sequence $U_{n}:{\mathcal{X}}^{n}\rightsquigarrow{\mathcal{Y}}_{n}$ and sequence of functions $f_{n}:{\mathcal{Y}}_{n}\to{\mathcal{Z}}_{n}$, we denote by $V=f(U)$ a channel sequence $V_{n}=f_{n}(U_{n}):{\mathcal{X}}^{n}\rightsquigarrow{\mathcal{Z}}_{n}$, $n\in{\mathbb{N}}$. Also, we denote by $\Phi_{n}=\log_{2}\max_{z\in{\mathcal{Z}}_{n}}\left|{f_{n}^{-1}(z)}\right|$.
@@
If $\Phi_{n}\in o(n)$, then ${\mathbf{C}}(U)={\mathbf{C}}(V)$. If, furthermore, at least one of ${\mathbf{I}}(U)$, ${\mathbf{I}}(V)$ exist, then they both exist and ${\mathbf{I}}(U)={\mathbf{I}}(V)$. Moreover, ${\mathbf{I}}{\mathbf{S}}(U)\iff{\mathbf{I}}{\mathbf{S}}(V)$.
@@
First, note that ${\mathbf{E}}_{p_{n}\circ V_{n}}{\left[{\mathbf{i}(V_{n},p_{n},x,y)}\right]}={\mathbf{E}}_{p_{n}\circ U_{n}}{\left[{\mathbf{i}(V_{n},p_{n},x,f(y))}\right]}$. Using Proposition 3 and the fact that $V_{n}$ is degraded with respect to $U_{n}$, for any sequence of input distributions $p$:
00 0 $\leq\lim_{n\to\infty}\frac{{\mathbf{I}}(U_{n},p_{n})-{\mathbf{I}}(V_{n},p_{n})}{n}={\mathbf{E}}_{p_{n}\circ U_{n}}{\left[{\frac{\mathbf{i}(U_{n},p_{n},x,y)-\mathbf{i}(V_{n},p_{n},x,f_{n}(y))}{n}}\right]}$
$\leq\lim_{n\to\infty}{\mathbf{E}}_{p_{n}\circ U_{n}}{\left[{\frac{\left|{\mathbf{i}(U_{n},p_{n},x,y)-\mathbf{i}(V_{n},p_{n},x,f_{n}(y))}\right|}{n}}\right]}\leq\lim_{n\to\infty}\frac{\Phi_{n}}{n}=0,$ (15)
which for optimal $p$ for $U$ implies that ${\mathbf{I}}(U)={\mathbf{I}}(V)$, if ether ${\mathbf{I}}(U)$ or ${\mathbf{I}}(V)$ exists.
Equation (15) implies that for any $p_{n}\in{\mathbb{D}}_{{\mathcal{X}}^{n}}$,
$\lim_{n\to\infty}{\mathbf{E}}_{p_{n}\circ U_{n}}{\left[{\frac{\left|{\mathbf{i}(U_{n},p_{n},x,y)-\mathbf{i}(V_{n},p_{n},x,f_{n}(y))}\right|}{n}}\right]}=0.$
Using the same technique as in Proposition 2, it can be easily shown that for any $\delta>0$ and $p_{n}\in{\mathbb{D}}_{{\mathcal{X}}^{n}}$,
$\lim_{n\to\infty}{\Pr_{p_{n}\circ U_{n}}{\left\{{\frac{\left|{\mathbf{i}(U_{n},p_{n},x,y)-\mathbf{i}(V_{n},p_{n},x,f_{n}(y))}\right|}{n}>\delta}\right\}}}=0.$ (16)
If ${\mathbf{I}}{\mathbf{S}}(U)$, then for arbitrary $\delta>0$, both (11) and (16) can be satisfied for $\delta/2$ simultaneously using some sequence of distributions $p^{*}$. The equality ${\mathbf{I}}(U)={\mathbf{I}}(V)$, proven above, and the triangle inequality imply
$\lim_{n\to\infty}{\Pr_{p^{*}_{n}\circ U_{n}}{\left\{{\left|{\frac{\mathbf{i}(V_{n},p^{*}_{n},x,f_{n}(y))}{n}-{\mathbf{I}}(V)}\right|>\delta}\right\}}}\leq\lim_{n\to\infty}{\Pr_{p^{*}_{n}\circ U_{n}}{\left\{{\frac{\left|{\mathbf{i}(U_{n},p^{*}_{n},x,y)-\mathbf{i}(V_{n},p^{*}_{n},x,f_{n}(y))}\right|}{n}>\frac{\delta}{2}}\right\}}}$
$+\lim_{n\to\infty}{\Pr_{p^{*}_{n}\circ U_{n}}{\left\{{\left|{\frac{\mathbf{i}(U_{n},p^{*}_{n},x,y)}{n}-{\mathbf{I}}(U)}\right|>\frac{\delta}{2}}\right\}}}=0.$ (17)
This implies ${\mathbf{I}}{\mathbf{S}}(V)$. The implication ${\mathbf{I}}{\mathbf{S}}(V)\implies{\mathbf{I}}{\mathbf{S}}(U)$ can be shown similarly.
@@
$\forall\varepsilon>0,\delta>0,\exists n^{*}_{0}:\forall n\geq n_{0},\exists p^{*}_{n}:~{}\Pr_{p^{*}\circ U_{n}(x,y)}{\left[{\frac{\mathbf{i}(U_{n},p^{*},x,y)}{n}>\alpha-\varepsilon}\right]}$ $>1-\delta.$ (18)
$\exists\varepsilon^{*}>0,\delta^{*}>0:\forall n_{0},\exists n^{*}\geq n_{0}:\forall p_{n},~{}\Pr_{p\circ U_{n^{*}}(x,y)}{\left[{\frac{\mathbf{i}(V_{n^{*}},p,x,f_{n^{*}}(y))}{n^{*}}\leq\alpha-\varepsilon^{*}}\right]}$ $\geq\delta^{*}.$ (19)
@@
$\exists\varepsilon^{*}>0,\delta^{*}>0:\begin{cases}\exists n^{*}_{0}:\forall n\geq n^{*}_{0},\exists p_{n}^{*}:~{}\Pr_{p_{n}^{*}\circ U_{n}(x,y)}{\left[{\mathbf{i}(U_{n},p_{n}^{*},x,y)/n>\alpha-\varepsilon^{*}/2}\right]}>1-\delta^{*}/2\\ \forall n_{0},\exists n^{*}\geq n_{0}:\forall p_{n^{*}},~{}\Pr_{p_{n^{*}}\circ U_{n^{*}}(x,y)}{\left[{\mathbf{i}(V_{n^{*}},p_{n^{*}},x,f_{n^{*}}(y))/{n^{*}}\leq\alpha-\varepsilon^{*}}\right]}\geq\delta^{*}\end{cases}$
So, the first inequality holds for all $n$ starting from $n^{*}_{0}$, and for each $n$ specific $p_{n}^{*}$ should be picked up. The second inequality holds for some $n$ larger than arbitrary $n_{0}$, for all $p_{n}$. Bringing these two conditions together, we can satisfy both inequalities as follows:
@@
  • since $n^{*}\geq n^{*}_{0}$, the first equality still holds for some specific selection of $p^{*}_{n^{*}}$. The second inequality will hold for $p_{n^{*}}^{*}$ as well, since it holds for any $p_{n^{*}}$.
@@
$\exists\varepsilon^{*}>0,\delta^{*}>0:\forall n_{0},\exists n\geq n_{0},\exists p_{n}:\begin{cases}\Pr_{p_{n}\circ U_{n}(x,y)}{\left[{\mathbf{i}(U_{n},p_{n},x,y)/n>\alpha-\varepsilon^{*}/2}\right]}>1-\delta^{*}/2\\ \Pr_{p_{n}\circ U_{n}(x,y)}{\left[{\mathbf{i}(V_{n},p_{n},x,f_{n}(y))/{n}\leq\alpha-\varepsilon^{*}}\right]}\geq\delta^{*}\end{cases}.$
Note that both probabilities are defined over the same probability space. The sum of the two probabilities is greater than $21+\delta^{*}/$ 2, which means that
${\Pr_{p_{n}\circ U_{n}(x,y)}{\left\{{\frac{\mathbf{i}{\left({U_{n},p_{n},x,y}\right)}-\mathbf{i}{\left({V_{n},p_{n},x,f_{n}(y)}\right)}}{n}>\frac{\varepsilon^{*}}{2}}\right\}}}>\frac{\delta^{*}}{2},$
which contradicts (16). ∎
Thus, one can apply such function sequences to a channel sequence $W$, changing neither ${\mathbf{I}}(W)$, nor ${\mathbf{C}}(W)$, nor information stability property.
@@
Consider channels $U:{\mathcal{X}}\rightsquigarrow{\mathcal{Y}}$ and $V:{\mathcal{X}}\rightsquigarrow{\mathcal{Z}}$, such that $V=f(U)$, $f:{\mathcal{Y}}\to{\mathcal{Z}}$. Consider partitions ${\mathcal{Y}}={\mathcal{A}}\sqcup{\mathcal{B}}$, ${\mathcal{Z}}={\mathcal{A}}^{\prime}\sqcup{\mathcal{B}}^{\prime}$, such that $f({\mathcal{A}})={\mathcal{A}}^{\prime}$ and $f$ is bijective between ${\mathcal{A}}$ and ${\mathcal{A}}^{\prime}$. Denote by $\beta(x)=\Pr_{U(y|x)}{\left[{y\in{\mathcal{B}}}\right]}$ the probability that the output of $U$ belongs to ${\mathcal{B}}$ given input $x\in{\mathcal{X}}$. If $\forall x\in{\mathcal{X}}:\beta(x)\leq\overline{\beta}$, then for any $p\in{\mathbb{D}}_{\mathcal{X}}$,
${\mathbf{I}}(U,p)-{\mathbf{I}}(V,p)\leq\overline{\beta}\log_{2}|{\mathcal{X}}|+\frac{1}{e\ln 2}.$ (20)
@@
By definition, ${\mathbf{I}}(U,p)={\mathbf{E}}_{p\circ U(x,y)}{\left[{\mathbf{i}(U,p,x,y)}\right]}$. The expectation can be expressed by two separate summations over ${\mathcal{A}}$ and ${\mathcal{B}}$:
@@
Let us analyze term $b$. First, introduce the modified distribution ${\widetilde{p}}(x)=p(x)\beta(x)/\Delta$, where $\Delta=\sum_{x}p(x)\beta(x)$. Note that $0\leq\Delta\leq\overline{\beta}$. Substituting $p(x)U(y|x)=\Delta{\widetilde{p}}(x)U^{\prime}(y|x)$ for $y\in{\mathcal{B}}$, one obtains
@@
$=\Delta\cdot{\mathbf{I}}(U^{\prime},{\widetilde{p}})+\Delta\sum_{y\in{\mathcal{B}},x\in{\mathcal{X}}}{\widetilde{p}}(x)U^{\prime}(y|x)\log_{2}\frac{\beta(x)}{\Delta}=\Delta{\left({{\mathbf{I}}(U^{\prime},{\widetilde{p}})-\log_{2}\Delta}\right)}+\sum_{y\in{\mathcal{B}},x\in{\mathcal{X}}}p(x)U^{\prime}(y|x)\beta(x)\log_{2}\beta(x)$
@@
Consider channel sequences $U_{n}:{\mathcal{X}}^{n}\rightsquigarrow{\mathcal{Y}}_{n}$ and $V_{n}:{\mathcal{X}}^{n}\rightsquigarrow{\mathcal{Z}}_{n}$, such that $V_{n}=f_{n}(U_{n})$, $f_{n}:{\mathcal{Y}}_{n}\to{\mathcal{Z}}_{n}$. Consider partitions ${\mathcal{Y}}_{n}={\mathcal{A}}_{n}\sqcup{\mathcal{B}}_{n}$, ${\mathcal{Z}}_{n}={\mathcal{A}}^{\prime}_{n}\sqcup{\mathcal{B}}^{\prime}_{n}$, such that $f_{n}({\mathcal{A}}_{n})={\mathcal{A}}^{\prime}_{n}$ and $f_{n}$ is bijective between ${\mathcal{A}}_{n}$ and ${\mathcal{A}}^{\prime}_{n}$. Denote by $\beta_{n}(x)=\Pr_{U_{n}(y|x)}{\left[{y\in{\mathcal{B}}_{n}}\right]}$, and $\overline{\beta}_{n}=\max_{x}\beta_{n}(x)$. If the output of channel $U_{n}$ almost always belongs to ${\mathcal{A}}_{n}$, i.e., $\lim_{n\to\infty}\overline{\beta}_{n}=0$, then ${\mathbf{C}}(U)={\mathbf{C}}(V)$. If, in addition, one of ${\mathbf{I}}(U)$, ${\mathbf{I}}(V)$ exists, then they both exist and ${\mathbf{I}}(U)={\mathbf{I}}(V)$. Moreover, ${\mathbf{I}}{\mathbf{S}}(U)\iff{\mathbf{I}}{\mathbf{S}}(V)$.
@@
If either ${\mathbf{I}}(U)$ or ${\mathbf{I}}(V)$ exists, then, using Proposition 5 and distributions $p^{*}_{n}$, optimal for $U_{n}$,
$0\leq{\mathbf{I}}(U)-{\mathbf{I}}(V)\leq\lim_{n\to\infty}\frac{{\mathbf{I}}(U_{n},p^{*}_{n})-{\mathbf{I}}(V_{n},p^{*}_{n})}{n}\leq\lim_{n\to\infty}{\left({\frac{\overline{\beta}_{n}n\log_{2}|{\mathcal{X}}|}{n}+\frac{1}{ne\ln 2}}\right)}=0,$
so ${\mathbf{I}}(U)={\mathbf{I}}(V)$.
Assume that ${\mathbf{I}}{\mathbf{S}}(U)$ is true. Note that, for any $p_{n}$,
$\lim_{n\to\infty}{\Pr_{p_{n}\circ U_{n}(x,y)}{\left\{{{\mathbf{i}}(U_{n},p_{n},x,y)\neq{\mathbf{i}}(V_{n},p_{n},x,f_{n}(y))}\right\}}}=\lim_{n\to\infty}{\Pr_{p_{n}\circ U_{n}(x,y)}{\left\{{y\in{\mathcal{B}}}\right\}}}\leq\lim_{n\to\infty}\overline{\beta}_{n}=0.$ (22)
Since ${\mathbf{I}}(U)={\mathbf{I}}(V)$, there exists a sequence of distributions $(p^{*}_{n})_{n\in{\mathbb{N}}}$, such that for any $\delta>0$:
$\lim_{n\to\infty}{\Pr_{p^{*}_{n}\circ V_{n}(x,z)}{\left\{{\left|{\frac{{\mathbf{i}}(V_{n},p^{*}_{n},x,z)}{n}-{\mathbf{I}}(V)}\right|>\delta}\right\}}}=\lim_{n\to\infty}{\Pr_{p^{*}_{n}\circ U_{n}(x,y)}{\left\{{\left|{\frac{{\mathbf{i}}(V_{n},p^{*}_{n},x,f_{n}(y))}{n}-{\mathbf{I}}(U)}\right|>\delta}\right\}}}$
$\leq\lim_{n\to\infty}\bigg{[}{\Pr_{p^{*}_{n}\circ U_{n}(x,y)}{\left\{{{\mathbf{i}}(V_{n},p^{*}_{n},x,f_{n}(y))\neq{\mathbf{i}}(U_{n},p^{*}_{n},x,y)}\right\}}}+{\Pr_{p^{*}_{n}\circ U_{n}(x,y)}{\left\{{\left|{\frac{{\mathbf{i}}(U_{n},p^{*}_{n},x,y)}{n}-{\mathbf{I}}(U)}\right|>\delta}\right\}}}\bigg{]},$
and by (22) and information stability of $U$, the limit is 00. Thus, ${\mathbf{I}}{\mathbf{S}}(U)\implies{\mathbf{I}}{\mathbf{S}}(V)$ is information stable. The other implication can be shown similarly.
According to (9), ${\mathbf{C}}(U)={\mathbf{C}}(V)$ can be proved by analyzing the probability of $\mathbf{i}(V_{n},p_{n},x,y)$ being less than $\alpha n$ for some value of $\alpha$. This can be done as follows:
@@
$\leq\Pr_{p_{n}\circ U_{n}(x,y)}{\left[{\mathbf{i}(V_{n},p_{n},x,f(y))\neq\mathbf{i}(U_{n},p_{n},x,y)}\right]}+\Pr_{p_{n}\circ U_{n}(x,y)}{\left[{\mathbf{i}(U_{n},p_{n},x,y)\leq\alpha n}\right]},$
where the first probability goes to zero according to (22). Thus,
@@
which means that the set of such $\alpha$’s for $U_{n}$ is a subset of such set for $V_{n}$, which implies ${\mathbf{C}}(V)\geq{\mathbf{C}}(U)$ and thus ${\mathbf{C}}(V)={\mathbf{C}}(U)$. ∎
If this proposition is combined with Proposition 4, the result is that one can apply any finite number of functions, which either have sufficiently small pre-image or sufficiently small probability of changing the output, without changing neither ${\mathbf{I}}$ nor ${\mathbf{C}}$ of the original channel sequence. The ${\mathbf{I}}{\mathbf{S}}$ property is also saved upon such transformations.
IV The Markov-IDS channel sequence
IV-A The IDS channel sequence
In [5] a general channel model, which represents any channel with independent insertions, deletions, and substitutions is considered. Consider a basic channel $W:{\mathcal{X}}\rightsquigarrow{\overline{\mathcal{Y}}}$, such that for each $x\in{\mathcal{X}}$ the expected length ${\mathbf{N}}(y)$ of the output sequence $y\in{\overline{\mathcal{Y}}}$ is finite: